OpenStudy (anonymous):

What volume of a 1.20M H2SO4 solution is needed for a reaction that requires 22.7g of solute?

OpenStudy (anonymous):

\[c =\frac{ n }{ v }\] c=concentration (measured in M) n=no. of moles (mol) v=volume (dm^3)

OpenStudy (anonymous):

can you carry on from there??

OpenStudy (anonymous):

So would it be 1.20=22.7/x? I'm not 100% sure how that would work out. Do I need to convert anything?

OpenStudy (anonymous):

you need to convert the mass to moles

OpenStudy (anonymous):

should i help with that??

OpenStudy (anonymous):

\[n = \frac{ m }{ M }\] n=no. of moles m=mass M=Molar mass (the same as mass number)

OpenStudy (anonymous):

you are welcome :)